Draw a sketch for a hand lever of effective length= 600 mm , hand force = 100 N

tensile stress = 7000 N/m^2 , shear sress = 6000 N/m^2


Determine ;

shaft diameter in the boss and in the bearing ,

boss diameter and length ,

width and thickness of arm .





1
Expert's answer
2021-03-11T06:55:57-0500

P x L = "\pi" /16 x q x d3

P - hand force; L - effective length; q - shear sress; d - shaft diameter in the boss

100 N x 0.6 m = "\pi" /16 x 6000 N/m2 x d3

d = 0.370 m = 370 mm 

The diameter of boss d2 = 1.6 x d = 1.6 x 370 mm = 592 mm

Lenght of boss l1 = (2 x P x L) / {t2 x "\sigma" x (d + t2)}

"\sigma" - tensile stress; t2 = thickness of boss = 0.3d = 0.3 x 0.37 m = 0.111 m

l1 = (2 x 100 N x 0.6 m) / {0.111 m x 7000 N/m2 x (0.370 m + 0.111 m)} = 0.321 m = 321 mm

P x (L2 + l2)1/2 = "\pi" /16 x q x d13

l = 2 x l1 = 2 x 0.321 m = 0.642 m; d1 - shaft diameter in the bearing

100 N x ((0.6 m)2 +(0.642 m)2)1/2 = "\pi" /16 x 6000 N/m2 x d13

d1 = 0.422 m = 422 mm